Solve:
log3(7x-1)-log3(x+1)=log3 5
Can you provide the steps to your answer

Answer by stanbon(75887)

log3(7x-1)-log3(x+1)=log3 5
----
Comment: Logs are exponents. The effect of adding
them is to multipl and the effect of subtracting
them is to divide.
------
Your Problem:
log3[(7x-1)/(x+1)] = log3(5)
===================
Taking the anti-log you get:
(7x-1)/(x+1) = 5
---------
Solve for "x":
6x-1 = 5x+5
x = 6
========================
Cheers,
Stan H.
========================